OpenStudy (anonymous):

toss a dice 1000 times how many times will you get a 6

OpenStudy (skullpatrol):

Nobody knows for sure :)

OpenStudy (skullpatrol):

toss a coin 1000 times how many times will you get a head?

OpenStudy (anonymous):

There is experimental probability and theoretical probability. Experimental probability means that if you toss a dice 1,000 times your outcome may vary. Theoretical probability means that there will be equally likely events. So in terms of theoretical probability you will get a 6 166.6 times (167).

OpenStudy (anonymous):

I found 167 by divided 1000 by 6, because there are six sides on a dice.

hartnn (hartnn):

probability of getting a 6 in one roll of a dice = 1/6 probability = number of favorable events N (getting a 6) / total number of events(=1000) = 1/6 so you just need N from N/1000 = 1/6